OpenStudy (anonymous):

if the Speedy family traveled 600 iles in 12 hours, with 1 hr stopped for dinner, 1/2hour for lunch and another 1/2 hr for restroom breaks, what wasthe rate of travel per hour that they were on the road? A. 45 mph B. 50 mph C. 55 mph D. 60 mph

OpenStudy (anonymous):

what is the total time they were on the road driving?

OpenStudy (anonymous):

12 i think

OpenStudy (anonymous):

no... remember they stopped a few times...

OpenStudy (anonymous):

then 14?

OpenStudy (anonymous):

how did u get 14?

OpenStudy (anonymous):

when they stopped for dinner, they are NOT driving.. so you need to subtract at least that much from the 12

OpenStudy (anonymous):

Ohhhhhh!

OpenStudy (texaschic101):

12 hours - 1 hour - 1/2 hour - 1/2 hour = 10 hours. They spent 10 hours actually driving.

OpenStudy (anonymous):

so 600 miles in 10 hours or \(\large \frac{600miles}{10hours}= \) ???

OpenStudy (anonymous):

thanks so it would be 60!?

OpenStudy (anonymous):

60 mph yes....

OpenStudy (texaschic101):

speed = distance divided by time (s = d/t)

OpenStudy (anonymous):

Thanks so much for the help (: